OpenStudy (anonymous):

HELP ME I GIVE RIBBON In which measurement situations could the most appropriate unit of measurement be millimeters? Choose all answers that are correct. A. length of a paper clip B. height of a person C. width of a finger D. length of a fishing boat

OpenStudy (anonymous):

@kathlyn98

OpenStudy (anonymous):

A

OpenStudy (anonymous):

Omg ty!

OpenStudy (anonymous):

wait I need 2 anwsers

OpenStudy (anonymous):

Answers*

OpenStudy (anonymous):

your welcome and c

OpenStudy (anonymous):

omg ty so much can u help me on another one?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

What is the perimeter of this pentagon? 

OpenStudy (anonymous):

I think is 50.1

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

yes you are right

OpenStudy (anonymous):

Ok ty

OpenStudy (anonymous):

Which equation could be used to find the perimeter of an isosceles triangle, with two sides 13 cm long, a base 10 cm long, and a height of 12 cm? A. P = (10 × 12) ÷ 2 B. P = 2 × 13 + 10 C. P = 13 + 10 + 12 D. P = 2 × 10 + 13

OpenStudy (anonymous):

A. = = (10 x 12) divided by 2. B. P = 2 X 13 + 10 C. P = 13 + 10 + 12 D. 2 X 10 + 13

OpenStudy (anonymous):

A. P = (10 x 12) divided by 2*

OpenStudy (anonymous):

it would be b

OpenStudy (anonymous):

oh k

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

 The perimeter of this shape is 108.5 in. What is the length of side d?

OpenStudy (anonymous):

82.6